# Toward A Causal Framework for Modeling Perception

### v3

@@ 3.1 The Representativeness Heuristic
Although in Definition 3.1 we care for the disagreement in probabilistic reasoning, we briefly discuss the mental heuristic commonly associated with perception as we view it in this work (Kahneman 2021). We do so to further support our focus on perception of probability. The representativeness heuristic is one among many heuristics humans employ that lead to bias in decision-making under uncertainty (Tversky and Kahneman 1974). Other heuristics, however, such as availability and anchoring bias, could be used to motivate our definition of perception.
We can view probabilities as a way to quantify how representative an instance $x$ is of a random variable $X$. Definition 3.1 coalesces into two individuals judging differently the representativeness of the instance $x\in X$. The representativeness heuristic is used in scenarios where we are asked to evaluate the degree to which one instance $x$ is representative of another instance $x^{\prime}$, leading us to evaluate poorly the representativeness of $x$ in terms of the probability $P(x)$. In such scenarios, we dwell into the question of resemblance between instances: the more one instance resembles another, the more representative it is of the other instance. The question “what is the probability that $x$ belongs to the class $X$?” asks “to what degree the instance $x$ resembles other known instances in the random variable $X$?” If the resemblance is high, then we judge the probability $P(x)$ that $x$ belongs to $X$ or, equivalently, that $x$ is generated by $X$ to be high. It helps to reason in terms of degree of representativeness, which translates naturally into how we understand probabilities.
@@
From Definition 3.2, it follows that if the known instance $x^{\prime}$ is viewed as representative of $X$, denoted, e.g., by a high-enough $P(x^{\prime})$ within the relevant context, then $x$ is also representative of $X$ as captured by $P(x)$. The degree of representativeness is clearly a function of what someone considers representative of $X$ as captured by $x^{\prime}$ and $P(x^{\prime})$. Receivers have their own $x^{\prime}$ that they use to construct $P(x)$.
@@
In Example 2, each admission officer resorts to their own degree of representativeness when evaluating the applicant with profile $\mathbf{x}=\langle x_{1}^{\prime},x_{2}^{\prime}\rangle$. The idea behind the competing $P_{R_{1}}(Y|x_{1},x_{2})$ and $P_{R_{2}}(Y|x_{1},x_{2})$ is that each officer compares the applicant relative to their own $\mathbf{x}^{\prime}$.
For our purposes, we view this process as representing individual experience. It is private information known only to each receiver. Both how such individual experience is constructed over time and elicited while reasoning by the receiver is beyond the scope of this work. We do, however, highlight another of Kahneman’s work: norm theory (Kahneman and Miller 1986), which theorizes how individuals respond to the familiarity of an event by recruiting and creating alternative scenarios. Under this theory, $R$ recruits a number of representations about an event. These representations are based on what $R$ views as a normal, with each scenario having a set of elements and each element having a set of features. These representations can be aggregated into a single scaled representation denoting the most common alternative, or the norm, according to $R$. Such a process could help formulate under a ML approach how individual experience is constructed, stored, and used by the receiver. We leave this for future work.
@@ 4 Formulating Causal Perception
We now extend Definition 3.1 to account for probabilistic causal reasoning. We are interested in capturing disagreement between receivers beyond the what is probability distribution by also capturing the what if probability distributions of the information sent by the sender. We rely on a SCM $\mathcal{M}$ to model the causal generative distribution $P(\mathbf{X})$. Notably, the single class case $P(X)$ is of no interest in this setting as it lacks a parent or child to reason about.
@@ 4.1 Why Causality?
Our modeling choice is based on three factors. First, under the premise that humans use probabilistic reasoning for decision-making under uncertainty (Kahneman 2011) as already argued in Section 3, the properties of a SCM $\mathcal{M}$ (Equation 1) offer a structured way to describe the information $\mathbf{X}$ as represented by $P(\mathbf{X})$. A SCM $\mathcal{M}$ allows to factorize $P(\mathbf{X})$ (Equation 3) and to reason under hypothetical scenarios about $P(\mathbf{X})$ (Equation 4), in principle, similar to how humans interpret information and accumulate knowledge (Woodward 2005; Pearl 2009; Pearl and Mackenzie 2018). Second, SCMs are increasingly used by ML researchers to approximate human-like reasoning (Schölkopf 2022; Schölkopf et al. 2021). Given this trend, our proposed framework would be compatible with the next wave of ML applications (e.g., van Steenkiste et al. (2019); Dittadi et al. (2021)). Third, SCMs, in particular through the DAG $\mathcal{G}$, are useful tools for engaging multiple stakeholders. For instance, DAGs have been used to draw the assumptions about $\mathbf{X}$ and its data generating model in sensitive settings like algorithmic discrimination testing (Álvarez and Ruggieri 2023) and synthetic data generation (Baumann et al. 2023). Here, the DAGs would be used to draw the individual experience of the receivers.
We do not view SCMs as equivalent to human reasoning. Rather, based on the above three factors, we simply view SCMs as a useful and pragmatic tool for formalizing human reasoning about $\mathbf{X}$ in terms of probabilities (similar, e.g., to Loftus (2024)). We are aware that this is not a view widely held within the fairness community (e.g., Hu and Kohler-Hausmann (2020)). Future work could explore non-causal approaches to extend Definition 3.1 that also capture reasoning about hypothetical scenarios.
@@ 4.2 Perception of Implied Probabilities
(Causal Perception) For receivers $R_{i},R_{j}\in\mathcal{R}$ with SCM $\mathcal{M}_{R_{i}}$ and $\mathcal{M}_{R_{i}}$ for the information $\mathbf{X}$ provided by sender $S\in\mathcal{S}$, causal perception occurs when, for a threshold ϵ+italic-$\epsilon\in\mathbb{R}^{+}$:
$\overline{d}\big{(}\mathcal{P}_{\mathbf{X}_{R_{i}}},\mathcal{P}_{\mathbf{X}_{R_{j}}}\big{)}>\epsilon$ (6)
where $\mathcal{P}_{\mathbf{X}_{R_{i}}}$ and $\mathcal{P}_{\mathbf{X}_{R_{j}}}$ represent the poset of all distributions of information $\mathbf{X}$ induced by $\mathcal{M}_{R_{i}}$ and $\mathcal{M}_{R_{i}}$, and $\overline{d}(\cdot,\cdot)$ denotes a suitable aggregated distance measure between two sets of probability distributions.
The Definition 4.1 extends Definition 3.1 into the realm of probabilistic causal reasoning. It defines perception beyond a disagreement on the representation of $\mathbf{X}$ in terms of probabilities by also accounting for any disagreement on the interventional distributions of the causal models. For instance, in Example 1 we would go from disagreeing on “what is the probability of an applicant being successful” to also disagreeing on “what is the probability of an applicant being successful had the applicant been different on one of its variables”. The latter kind of disagreement is inherently linked to the SCM of each receiver.
For Equation (6) we propose in practice to compute the distance as an aggregation of these distributions, such as an average or a maximum (see, e.g., Goldenberg and Webb (2019)). Formally, given a set of interventions $\mathcal{I}_{\mathbf{X}}$, we consider the aggregated perception for $R_{i}$ and $R_{j}$ in $\mathcal{R}$ as:
@@
such that $P^{do(l)}_{R_{i}}(\mathbf{X})$ denotes the interventional distribution in the SCM $\mathcal{M}_{R_{i}}$ of receiver $R_{i}$ after performing an intervention $l$. The same applies for receiver $R_{j}$ The choice of the distance function $d(\cdot,\cdot)$ and the aggregation function $\Phi$ will depend on the context, meaning the kind of disagreement we wish to capture. We could, for instance, define $d(\cdot,\cdot)$ as the Kullback–Leibler divergence and $\Phi$ as the average.
@@
In the case of a specific intervention, or $\mathcal{I}_{\mathbf{X}}=\{l\}$, Definition 4.1 becomes:
$d\big{(}P_{R_{i}}(X)^{do(l)},P_{R_{j}}(X)^{do(l)}\big{)}>\epsilon.$ (8)
The above equation highlights how Definitions 3.1 and 4.1 relate to each other. Here, we study the disagreement of the interventional distribution resulting from the l intervention. We can generate these interventional distributions, respectively, via $\mathcal{M}_{R_{i}}$ and $\mathcal{M}_{R_{j}}$.
@@ 4.3 Structural and Parametrical Perception
Central to Definition 4.1 are the SCMs of each receiver for $P(\mathbf{X})$. Importantly, if we assume that receivers must draw all variables $\mathbf{X}$, then two kinds of causal perception emerge: structural and parametrical. This assumption simply implies that the receivers, who receive the same information $\mathbf{X}$ from the sender, must use all elements in $\mathbf{X}$, meaning each $X\in\mathbf{X}$ is also a variable in $\mathcal{M}$. Recall that we consider the particular case of ANM (2) for illustrative purposes.
@@
(Structural Causal Perception) It occurs when receivers $R_{j},R_{j}\in\mathcal{R}$ disagree on the cause-effect pairs in, respectively, $\mathcal{M}_{R_{i}}$ and $\mathcal{M}_{R_{j}}$ for $P(\mathbf{X})$. It implies distinct DAGs $\mathcal{G}_{R_{i}}$ and $\mathcal{G}_{R_{j}}$.
@@
(Parametrical Causal Perception) It occurs when receivers $R_{j},R_{j}\in\mathcal{R}$ agree on the cause-effect pairs for $P(\mathbf{X})$, meaning the their DAGs $\mathcal{G}_{R_{i}}$ and $\mathcal{G}_{R_{j}}$ are the same, but disagree on the sets of structural equations $\mathbf{F}$ in respectively, $\mathcal{M}_{R_{i}}$ and $\mathcal{M}_{R_{j}}$.
Suppose, for instance, that $\mathbf{X}=\{X_{1},X_{2}\}$. In structural causal perception, $R_{i}$ draws $X_{1}\rightarrow X_{2}$ while $R_{j}$ draws $X_{2}\rightarrow X_{1}$. The distinct DAGs imply structurally different equations in the SCM as cause-effect pairs are reversed. In parametrical causal perception, assume $R_{i}$ and $R_{j}$ agree on $X_{1}\rightarrow X_{2}$. Broadly, it means both will have a structural equation of the form $X_{2}:=\beta\times X_{1}+U$ in their SCM. However, for $R_{i}$ $\beta>0$ while for $R_{j}$ $\beta<0$. These two equations, thus, are parametrized differently coefficient-wise.
@@
Let us consider scenario (ii) from Example 1 and admissions officers $R_{1}$ and $R_{2}$ for choosing which of the two applicants are admitted. Note that implied by the same score from $f$ is that these two applicants have a comparable profile $\mathbf{x}=\langle x_{1},x_{2}\rangle$. Suppose that the officers resort to the applicants zip code $Z$, meaning both reason about $P(Y|X_{1},X_{2},Z)$, by asking what would have been the score of the applicant had she or he been from the $Z=z$ zip code? We implement such hypothetical question using the $do(Z:=z)$ on the SCMs $\mathcal{M}_{R_{1}}$ and $\mathcal{M}_{R_{2}}$ and, thus, looking at the interventional joint probability distribution that is $P(Y|X_{1},X_{2},do(Z:=z))$.
First, let us consider the case of structural perception. Let $G_{R_{1}}=\{Z\rightarrow X_{1},Z\rightarrow X_{2},X_{2}\rightarrow X_{1},X_{1}\rightarrow Y,X_{2}\rightarrow Y\}$ and $G_{R_{2}}=\{Z\rightarrow X_{2},X_{2}\rightarrow X_{1},X_{1}\rightarrow Y,X_{2}\rightarrow Y\}$, meaning $R_{2}$ does not agree with $R_{1}$ that $Z\rightarrow X_{1}$. Given these DAGs, each receiver factorizes the joint probability distribution of interest differently when answering this what if question. We have $P(Y|X_{1},X_{2})P(X_{2}|Z)P(X_{1}|X_{2},Z)P(Z)$ under $\mathcal{G}_{R_{1}}$ and $P(Y|X_{1},X_{2})P(X_{2}|Z)P(X_{1}|X_{2})P(Z)$ under $\mathcal{G}_{R_{2}}$. $\mathcal{G}_{R_{1}}$ corresponds to the following set of structural equations:
@@
and $\mathcal{G}_{R_{2}}$ corresponds to a similar set of structural equations with the difference that (10) becomes:
@@
as we must exclude the term $\beta_{2}\cdot Z$ since there is no $Z\rightarrow X_{1}$ in $\mathcal{G}_{R_{2}}$. In reasoning about $Z$ and its impact on the other variables, each receiver considers distinct interventional distributions due to the different factorizations of $P(\mathbf{X})$. $R_{1}$ constructs $P(Y|X_{1},X_{2})P(X_{2}|z)P(X_{1}|X_{2},z)P(z)$ while $R_{2}$ $P(Y|X_{1},X_{2})P(X_{2}|z)P(X_{1}|X_{2})P(z)$. In terms of the set if structural equations, we substitute each $Z$ for the value $z$. Thus, computing the hypothetical $G$ for $R_{1}$:
@@
where, ultimately, the term $\beta_{4}\cdot\beta_{2}\cdot z$ is missing for $R_{2}$’s $G$ score w.r.t. $R_{1}$. Such a difference could determine which candidate each officer chooses based on $Z$.
Now, let us consider the case of parametrical perception. Let $R_{1}$ and $R_{2}$ agree on $\mathcal{G}_{R_{1}}$. However, for equation (10), suppose $R_{1}$ considers $\alpha_{1}$ and $R_{2}$ considers $\alpha_{2}$ as the causal weight of $Z\rightarrow X_{1}$ (what is now $\beta_{2}$) such that $\alpha_{1}\gg\alpha_{2}$. Here, the same reasoning applies as with structural causal perception when constructing $P(Y|X_{1},X_{2},do(Z:=z))$. The difference is that both receivers arrive at (14) after intervening $Z$ with a difference in the $\beta_{4}\cdot\beta_{2}\cdot z$ term. This term is $\beta_{4}\cdot\alpha_{1}\cdot z$ for $R_{1}$ and $\beta_{4}\cdot\alpha_{2}\cdot z$ for $R_{2}$, where $\beta_{4}\cdot\alpha_{1}\cdot z>\beta_{4}\cdot\alpha_{2}\cdot z$. Such a difference once again could determine which candidate each officer chooses based on $Z$.
Overall, note that the term in question in both kinds of causal perception, represent a potentially lower or higher SAT score because of the applicants zip code (read, neighborhood composition) that would then enter the classifier $f$. We do this on purpose as the link between wealth (captured by $Z$) and aptitude (captured by $X_{1}$) is a devise one and of interest to human decision-makers (Kleinberg et al. 2019).
The central idea with these two kinds of causal perception is that they account for why the two receivers in Definition 4.1 disagree on the poset of implied distributions of SCM $\mathcal{M}$. The disagreement, by definition, comes from each receiver’s SCM. Intuitively, with Definition 4.2 the disagreement stems from different graphical structures, and with Definition 4.3 from different model parametrizations.
We view the SCMs as the individual experience of each receiver. We, once again, take for granted the construction and elicitation of these individual experiences. However, while Definition 3.1 could be explored deeper later on via, e.g., the representativeness heuristic, there is a lack for an equivalent framework to support Definition 4.1. In the next section, we propose a causal modeling framework that motivates a receiver’s individual experience as a SCM.

### v4

@@ 3.1 The Representativeness Heuristic
Although Definition 3.1 cares only for the disagreement in probabilistic reasoning, we briefly discuss the mental heuristic commonly associated with perception as we view it in this work. We do so to further support our focus on perception of probability. The representativeness heuristic is one among many heuristics used by humans that lead to biased decision-making under uncertainty (Tversky and Kahneman 1974). Other heuristics, however, such as availability and anchoring, could also motivate our definition of perception.
Probabilities can be understood as quantifying how representative an instance $x$ is of a random variable $X$. Under this premise, for instance, Definition 3.1 reflects the setting in which two individuals can judge differently the representativeness of the same instance $x$ relative to the random variable $X$. Similarly, the representativeness heuristic is used when humans assess how much one instance $x$ resembles another instance $x^{\prime}$ that is known to belong to $X$, which may distort the estimation of $P(x)$. This is because, in such a case, reasoning centers on resemblance between instances. The more one instance resembles the other instance, the more representative the former is of the latter instance. The question “What is the probability that $x$ belongs to the random variable XX?” becomes “To what degree the instance $x$ resembles the other known instance $x^{\prime}$ in the random variable XX?”. If the resemblance is high, then we judge the probability P(x)P(x)—read as $x$ belongs to $X$ or, equivalently, $x$ is generated by XX—to be high. Here, it helps overall to reason in terms of degree of representativeness, which translates naturally into how we understand probabilities.
@@
It follows from Definition 3.2 that, if the known instance $x^{\prime}$ is viewed as representative of $X$, denoted by a high-enough $P(x^{\prime})$ within the relevant context, then $x$ is also representative of $X$ as captured by $P(x)$. The degree of representativeness is clearly a function of what someone considers representative of $X$ as captured by $x^{\prime}$ and $P(x^{\prime})$. With respect to our framework for perception, it implies that receivers have their own $x^{\prime}$ that they use to construct $P(x)$.
@@
The occurrence of perception in Example 2 is driven by each admissions officer resorting to their own degree of representativeness when evaluating the deferred applicant with profile $\mathbf{x}=\langle x_{1},x_{2}\rangle$. Underlying the competing $P_{R_{1}}(Y|x_{1},x_{2})$ and $P_{R_{2}}(Y|x_{1},x_{2})$ is that each officer compares $\mathbf{x}$ relative to their own $\mathbf{x}^{\prime}$.
For our purposes, we view this process as representing individual experience. It is private information known only to the receivers. How such individual experience is constructed over time and elicited while reasoning by the receiver is beyond the scope of this work. We do, however, highlight another of Kahneman’s work: norm theory (Kahneman and Miller 1986), which theorizes how individuals respond to the familiarity of an event by recruiting and creating alternative scenarios. Under this theory, $R$ recruits a number of representations about an event. These representations are based on what $R$ views as a normal, with each scenario having a set of elements and each element having a set of features. These representations can be aggregated into a single scaled representation denoting the most common alternative, or the norm, according to $R$. Such a process could help formulate under a ML approach how individual experience is constructed, stored, and used by the receiver. We leave this for future work.
@@ 4 Formulating Causal Perception
We now extend Definition 3.1 to account for probabilistic causal reasoning. We want to capture disagreement between receivers beyond the what is probability distribution by also considering the what if probability distributions of the information sent by the sender. We rely on a SCM $\mathcal{M}$ to model the causal generative distribution $P(\mathbf{X})$. Notably, the single random variable case $P(X)$ is of no interest here as it lacks a parent or child to reason about.
@@ 4.1 Why Causality?
Our modeling choice is based on three factors. First, under the premise that humans use probabilistic reasoning for decision-making under uncertainty (Kahneman 2011) as argued in Section 3, the properties of a SCM $\mathcal{M}$ (Equation 1) offer a structured way to describe the information $\mathbf{X}$ as represented by $P(\mathbf{X})$. A SCM $\mathcal{M}$ allows to factorize $P(\mathbf{X})$ (Equation 3) and to reason under hypothetical scenarios about $P(\mathbf{X})$ (Equation 4), in principle, similar to how humans interpret information and accumulate knowledge (Woodward 2005; Pearl 2009; Pearl and Mackenzie 2018). Second, SCMs are increasingly used by ML researchers to approximate human-like reasoning (Schölkopf 2022; Schölkopf et al. 2021). Given this trend, our proposed framework is compatible with the next wave of ML applications (see, e.g., van Steenkiste et al. (2019); Dittadi et al. (2021)). Third, a SCM, in particular through its DAG $\mathcal{G}$, is a useful tools for engaging multiple stakeholders. A DAG, for instance, can be used to draw the assumptions about $\mathbf{X}$ and its data generating model in sensitive settings like discrimination testing (Álvarez and Ruggieri 2023, 2025) and synthetic data generation (Baumann et al. 2023). The DAG can essentially “draw” relevant individual experience.
We do not view SCMs as equivalent to human reasoning. Rather, based on these three factors, we simply view SCMs as a useful and pragmatic tool for formalizing human reasoning about $\mathbf{X}$ in terms of probabilities (similar, e.g., to Loftus (2024)). We are aware that this is not a view widely held within the fairness community (see, e.g., Hu and Kohler-Hausmann (2020)). Future work could explore non-causal approaches to extend Definition 3.1 that also capture reasoning about hypothetical scenarios.
@@ 4.2 Perception of Implied Probabilities
(Causal Perception) For receivers $R_{i},R_{j}\in\mathcal{R}$ with SCM $\mathcal{M}_{R_{i}}$ and $\mathcal{M}_{R_{j}}$ for the information $\mathbf{X}$ provided by sender $S\in\mathcal{S}$, causal perception occurs when for a threshold $\epsilon\in\mathbb{R}^{+}$:
$\overline{d}\left(\mathcal{P}_{\mathbf{X}_{R_{i}}},\mathcal{P}_{\mathbf{X}_{R_{j}}}\right)>\epsilon$ (6)
where $\mathcal{P}_{\mathbf{X}_{R_{i}}}$ and $\mathcal{P}_{\mathbf{X}_{R_{j}}}$ represent the poset of all distributions of information $\mathbf{X}$ induced by $\mathcal{M}_{R_{i}}$ and $\mathcal{M}_{R_{j}}$, and $\overline{d}(\cdot,\cdot)$ denotes a suitable aggregated distance measure between two sets of probability distributions.
The Definition 4.1 extends Definition 3.1 into the realm of probabilistic causal reasoning. It defines perception beyond a disagreement on the representation of $\mathbf{X}$ in terms of probabilities by also accounting for any disagreement on the interventional distributions of the causal models. In Example 1, for instance, we would go from disagreeing on “What is the probability of an applicant being successful?” to also disagreeing on “What is the probability of an applicant being successful had the applicant been different on one of its variables?”. The latter kind of disagreement is inherently linked to the SCM of each receiver.
For Equation (6), we propose in practice to compute the distance as an aggregation of these distributions, such as an average or a maximum (see, e.g., Goldenberg and Webb (2019)). Formally, given a set of interventions $\mathcal{I}_{\mathbf{X}}$, we consider the aggregated perception for $R_{i}$ and $R_{j}$ in $\mathcal{R}$ as:
@@
such that $P^{do(l)}_{R_{i}}(\mathbf{X})$ denotes the l-interventional distribution in the SCM $\mathcal{M}_{R_{i}}$ of receiver $R_{i}$ after performing the l-intervention in $\mathcal{I}_{\mathbf{X}}$. The same applies for receiver $R_{j}$ and its SCM $\mathcal{M}_{R_{j}}$. The choice of the distance function $d(\cdot,\cdot)$ and the aggregation function $\Phi$ in Equation (7) will depend on the context and the kind of disagreement we wish to capture. We could, for instance, define $d(\cdot,\cdot)$ as the Kullback–Leibler divergence and $\Phi$ as the average.
@@
Following up the previous remark, in the case of a specific l-intervention, or $\mathcal{I}_{\mathbf{X}}=\{l\}$, Definition 4.1 becomes:
$d\left(P_{R_{i}}(X)^{do(l)},P_{R_{j}}(X)^{do(l)}\right)>\epsilon.$ (8)
Equation (8) highlights how Definitions 3.1 and 4.1 relate to each other as the difference in, respectively, the observational and the l-interventional distributions. Here, we study the disagreement of the interventional distribution resulting from the l-intervention. We can generate these interventional distributions via the SCMs $\mathcal{M}_{R_{i}}$ and $\mathcal{M}_{R_{j}}$.
@@ 4.3 Structural and Parametrical Perception
Central to Definition 4.1 are the SCMs of each receiver for $P(\mathbf{X})$. If we assume that receivers draw all variables $\mathbf{X}$ using a SCM $\mathcal{M}$, then two kinds of causal perception emerge: structural and parametrical. This assumption implies that the receivers, who receive the same information $\mathbf{X}$ from the sender, use all elements in $\mathbf{X}$, meaning each $X\in\mathbf{X}$ is also a variable in the SCM $\mathcal{M}$. We consider the particular case of ANM (Equation 2) for illustrative purposes.
@@
(Structural Causal Perception) It occurs when receivers $R_{i},R_{j}\in\mathcal{R}$ disagree on the cause-effect pairs in, respectively, $\mathcal{M}_{R_{i}}$ and $\mathcal{M}_{R_{j}}$ for $P(\mathbf{X})$. It implies distinct DAGs $\mathcal{G}_{R_{i}}$ and $\mathcal{G}_{R_{j}}$.
@@
(Parametrical Causal Perception) It occurs when receivers $R_{i},R_{j}\in\mathcal{R}$ agree on the cause-effect pairs for $P(\mathbf{X})$, meaning the their DAGs $\mathcal{G}_{R_{i}}$ and $\mathcal{G}_{R_{j}}$ are the same, but disagree on the sets of structural equations $\mathbf{F}$ in respectively, $\mathcal{M}_{R_{i}}$ and $\mathcal{M}_{R_{j}}$.
For instance, let $\mathbf{X}=\{X_{1},X_{2}\}$. In structural causal perception, $R_{i}$ draws $X_{1}\rightarrow X_{2}$ while $R_{j}$ draws $X_{2}\rightarrow X_{1}$. The distinct DAGs imply structurally different equations in the SCM as cause-effect pairs are reversed. In parametrical causal perception, assume $R_{i}$ and $R_{j}$ agree on $X_{1}\rightarrow X_{2}$. Broadly, it means both will have a structural equation of the form $X_{2}:=\beta\times X_{1}+U$ in their SCM. However, for $R_{i}$ $\beta>0$ while for $R_{j}$ $\beta<0$. These two equations are parametrized differently coefficient-wise.
@@
Let us consider Example 1, scenario (ii). Suppose two admissions officers, $R_{1}$ and $R_{2}$, are tasked to choose independently which of the two applicants tied by the score $G$ from the ML model $f$ is admitted. Two additional points to consider. First, the fact that $f$ gives the same score to both applicants implies comparable profiles, which can be summarized by a single $\mathbf{x}=\langle x_{1},x_{2}\rangle$. These two applicants are indistinguishable to $f$. Second, this fact allows the officers to derive the decision $Y$ without $G$. They can now consider other aspects of the applicants’ profiles. Further suppose the officers resort to the applicants’ zip code $Z$ for the tiebreaker. In particular, both officers reason about $P(Y|X_{1},X_{2},Z)$ by asking “What would have been the score of an applicant had she or he been from the $Z=z$ zip code?” We can represent such hypothetical question using the $do(Z:=z)$ on the SCMs $\mathcal{M}_{R_{1}}$ and $\mathcal{M}_{R_{2}}$ of each officer, thus, looking at the interventional joint probability distribution that is $P(Y|X_{1},X_{2},do(Z:=z))$.
Let us first consider the case of structural perception. Suppose two competing DAGs for $\{Y,X_{1},X_{2},Z\}$: let $G_{R_{1}}=\{Z\rightarrow X_{1},Z\rightarrow X_{2},X_{2}\rightarrow X_{1},X_{1}\rightarrow Y,X_{2}\rightarrow Y\}$ and $G_{R_{2}}=\{Z\rightarrow X_{2},X_{2}\rightarrow X_{1},X_{1}\rightarrow Y,X_{2}\rightarrow Y\}$, meaning $R_{2}$ does not agree with $R_{1}$ that $Z\rightarrow X_{1}$. Given these DAGs, each office factorizes the interventional joint probability distribution of interest differently when answering the what if question. We have $P(Y|X_{1},X_{2})P(X_{2}|Z)P(X_{1}|X_{2},Z)P(Z)$ under $\mathcal{G}_{R_{1}}$ and $P(Y|X_{1},X_{2})P(X_{2}|Z)P(X_{1}|X_{2})P(Z)$ under $\mathcal{G}_{R_{2}}$. The following set of structural equations corresponds to $\mathcal{G}_{R_{1}}$:
@@
and a similar set of structural equations corresponds to $\mathcal{G}_{R_{2}}$ with the key difference that (10) becomes:
@@
as we exclude the term $\beta_{2}\cdot Z$ since there is no $Z\rightarrow X_{1}$ cause-effect pair in $R_{2}$’s DAG. When reasoning about $Z$ and its impact on the other variables, each officer considers distinct interventional distributions due to the different factorizations of $P(\mathbf{X})$ as well as the competing set of structural equations. $R_{1}$ constructs $P(Y|X_{1},X_{2})P(X_{2}|z)P(X_{1}|X_{2},z)P(z)$ while $R_{2}$ $P(Y|X_{1},X_{2})P(X_{2}|z)P(X_{1}|X_{2})P(z)$. Similarly, in terms of the set of structural equations, when substituting each $Z$ for the value $z$ using the do-operator for $R_{1}$ we obtain:
@@
where the term $\beta_{4}\cdot\beta_{2}\cdot z$ is missing for $R_{2}$ with respect to $R_{1}$ when making the decision. Such a difference can lead to a different applicant chosen by each officer based on $Z$.
Let us now consider the case of parametrical perception. Suppose $R_{1}$ and $R_{2}$ agree on the previous DAG $\mathcal{G}_{R_{1}}$. However, for equation (10) suppose $R_{1}$ considers $\alpha_{1}$ and $R_{2}$ considers $\alpha_{2}$ as the causal weight of $Z\rightarrow X_{1}$ (what is now $\beta_{2}$) such that $\alpha_{1}\gg\alpha_{2}$. The same reasoning applies as with causal perception when constructing the interventional distribution $P(Y|X_{1},X_{2},do(Z:=z))$. The difference is that both receivers arrive at (14) after intervening $Z$ with but each has a distinct $\beta_{4}\cdot\beta_{2}\cdot z$ term. This term is $\beta_{4}\cdot\alpha_{1}\cdot z$ for $R_{1}$ and $\beta_{4}\cdot\alpha_{2}\cdot z$ for $R_{2}$, where $\beta_{4}\cdot\alpha_{1}\cdot z>\beta_{4}\cdot\alpha_{2}\cdot z$. Such a difference once again can determine which candidate each officer chooses based on $Z$.
The term in question in both kinds of causal perception represents a potentially lower or higher SAT score because of the applicants zip code (read, neighborhood composition). We do this on purpose as the link between wealth (captured by $Z$) and aptitude (captured by $X_{1}$) is a devise one and of interest (Kleinberg et al. 2019).
The central idea with the two kinds of causal perception is that they account for why the two receivers in Definition 4.1 disagree on the poset of implied distributions from a SCM $\mathcal{M}$. The disagreement, by definition, comes from each receiver’s SCM: with Definition 4.2 the disagreement stems from different graphical structures while with Definition 4.3 from different model parametrizations. These competing SCMs represent the individual experience of each receiver. Here we take for granted the construction and elicitation of these individual experiences. Notably, though, while in Definition 3.1 we made use of the representativeness heuristic to better understand individual experience in a probabilistic way, there is a lack for an equivalent framework to support Definition 4.1. Therefore, in the next section we propose a causal modeling framework that motivates a receiver’s individual experience as a SCM $\mathcal{M}$.